### v3

# Properties of resonance graphs that are daisy cubes

Abstract
Let $G$ be a plane elementary bipartite graph whose infinite face is forcing. We show that there is a bijection between the set of maximal resonant sets of $G$ and the set of maximal hypercubes of its resonance graph $R(G)$. Moreover, if $G$ is a peripherally 2-colorable graph whose inner dual is $G^{*}$, we then establish a bijection between the set of maximal hypercubes of its resonance graph $R(G)$ and the set of maximal independent sets of $G^{*}$. Next, we present a characterization on when the resonance graph of a plane elementary bipartite graph $G$ is a daisy cube in terms of the daisy cube constructed from the set of all independent sets of $G^{*}$. The obtained result is also extended to plane weakly elementary bipartite graphs. Finally, two algorithms which provide a proper labelling for the vertex set of $R(G)$ as a daisy cube are presented.
keywords: daisy cube, independent set, peripherally 2-colorable graph, plane (weakly) elementary bipartite graph, resonance graph
Simon Brezovnika,b, Zhongyuan Chec, Niko Tratnikd,b,
Petra Žigert Pleteršeke,d
a University of Ljubljana, Faculty of Mechanical Engineering, Slovenia
b Institute of Mathematics, Physics and Mechanics, Ljubljana, Slovenia
c Department of Mathematics, Penn State University, Beaver Campus, Monaca, USA
d University of Maribor, Faculty of Natural Sciences and Mathematics, Slovenia
e University of Maribor, Faculty of Chemistry and Chemical Engineering, Slovenia
simon.brezovnik@fs.uni-lj.si, zxc10@psu.edu,
niko.tratnik@um.si, petra.zigert@um.si
1 Introduction
Resonance graphs, also termed as $Z$-transformation graphs, depict the relationships among perfect matchings, also known as Kekulé structures in chemistry. Originating separately from the works of chemists [11, 9, 10] and mathematicians [24], these graphs were initially focused on benzenoid graphs, and later expanded to include plane bipartite graphs [14, 25, 28, 29]. This extension broadened the applicability of the concept. Research on resonance graphs of benzenoid graphs involving hypercubes is documented in several sources such as [15, 16, 18], while some structural aspects of resonance graphs of plane bipartite graphs are summarized in an early survey paper [22].
Daisy cubes were introduced in [12] as a subfamily of partial cubes which contains Fibonacci cubes and Lucas cubes. A binary coding of length $n$ for the vertex set of a graph $G$ isomorphic to a daisy cube is called a proper labelling if it induces an isometric embedding of $G$ into the hypercube $Q_{n}$ as a daisy cube [17]. Recent advancements, including characterizations and embeddings of daisy cubes, were discussed in [17, 20].
The linkage between resonance graphs and daisy cubes was initially explored in [30], revealing that the resonance graph of a kinky benzenoid graph is a daisy cube. Moreover, characterizations of catacondensed even ring systems and 2-connected outerplane bipartite graphs whose resonance graphs are daisy cubes were provided in [2] and [3], respectively. We further characterized plane bipartite graphs whose resonance graphs are daisy cubes in terms of peripherally 2-colorable graphs in [1]: the resonance graph of a plane bipartite graph $G$ is a daisy cube if and only if $G$ is a plane weakly elementary bipartite graph whose elementary components different from $K_{2}$ are peripherally 2-colorable.
This paper builds upon and extends the results from [1]. We first show that for a plane elementary bipartite graph whose infinite face is forcing, there is a bijection between the set of maximal resonant sets of such a graph and the set of maximal hypercubes of its resonance graph. Peripherally 2-colorable graphs are a special type of plane elementary bipartite graphs whose infinite face is forcing. For a peripherally 2-colorable graph $G$, we further show that there is a bijection between the set of maximal hypercubes of its resonance graph $R(G)$ and the set of maximal independent sets of its inner dual $G^{*}$. We then construct a daisy cube from the set of all independent sets of any graph, and present a characterization on when the resonance graph of a plane weakly elementary bipartite graph $G$ is a daisy cube in terms of the daisy cube generated by the set of all independent sets of the inner dual of the subgraph of $G$ obtained by removing all forbidden edges.
Finally, we write two algorithms which provide a proper labelling for all perfect matchings of a peripherally 2-colorable graph $G$ forming the vertex set of $R(G)$ as a daisy cube.
2 Preliminaries
2.1 Daisy cubes
Let $G$ be a graph with the vertex set $V(G)$ and the edge set $E(G)$. A set of pairwise nonadjacent vertices of $G$ is called an independent set. Moreover, a maximal independent set of $G$ is an independent set that is not a proper subset of any other independent set of $G$. The disjoint union of graphs $G_{1}$ and $G_{2}$, denoted as $G_{1}\cup G_{2}$, is the graph whose vertex set is a disjoint union of $V(G_{1})$ and $V(G_{2})$, and edge set is a disjoint union of $E(G_{1})$ and $E(G_{2})$. A subgraph of $G$ induced by a vertex subset $X\subseteq V(G)$ is denoted as $\langle X\rangle$. If $G$ is a graph whose vertex set is a poset $(V(G),\leq)$, then the maximal elements of the poset $(V(G),\leq)$ are called the maximal vertices of $G$, and the minimum element (if exists) of the poset $(V(G),\leq)$ is called the minimum vertex of $G$.
A Cartesian product $\Box_{i=1}^{n}G_{i}$ of $n$ graphs $G_{1},G_{2},\ldots G_{n}$, where $n\geq 2$, is a graph whose vertex set is $V(G_{1})\times V(G_{2})\times\cdots\times V(G_{n})$ and two vertices $(x_{1},\ldots,x_{n})$ and $(y_{1},\ldots,y_{n})$ are adjacent in $\Box_{i=1}^{n}G_{i}$ if there exists an $i\in\{1,\ldots,n\}$ such that $x_{i}y_{i}$ is an edge of $G_{i}$, and $x_{j}=y_{j}$ for any $j\in\{1,\ldots,n\}\setminus\{i\}$. An $n$-dimensional hypercube $Q_{n}$ is a Cartesian product $\Box_{i=1}^{n}K_{2}$, while $Q_{0}$ is a one-vertex graph, and $Q_{1}$ is the one-edge graph. We say that a hypercube is nontrivial if is not a one-vertex graph.
An induced subgraph $H$ of $G$ is an isometric subgraph if $d_{H}(u,v)=d_{G}(u,v)$ for every two vertices $u$ and $v$ of $H$, that is, the distance between any two vertices in $H$ is equal to their distance in $G$. Partial cubes are isometric subgraphs of hypercubes. Djoković-Winkler relation (briefly, relation $\Theta$) plays an important role in the study of partial cubes. Two edges $x_{1}x_{2}$ and $y_{1}y_{2}$ of a connected graph $G$ are in relation $\Theta$ if $d_{G}(x_{1},y_{1})+d_{G}(x_{2},y_{2})\neq d_{G}(x_{1},y_{2})+d_{G}(x_{2},y_{1})$. The relation $\Theta$ is an equivalence relation on the edge set of any partial cube, see [8].
Let integer $n\geq 1$. Let $(\mathcal{B}^{n},\leq)$ be a poset on the set of all binary strings of length $n$ with the partial order $u_{1}u_{2}\ldots u_{n}\leq v_{1}v_{2}\ldots v_{n}$ if $u_{i}\leq v_{i}$ for all $1\leq i\leq n$. A daisy cube generated by $X\subseteq\mathcal{B}^{n}$ is an induced subgraph of $Q_{n}$, and defined as $Q_{n}(X)=\langle\{u\in\mathcal{B}^{n}\ |\ u\leq x\textrm{ for some }x\in X\}\rangle$.
Moreover, if $\widehat{X}$ is the anti-chain consisting of the maximal elements of the poset $(X,\leq)$, then $Q_{n}(X)=Q_{n}(\widehat{X})$, see [12]. Hence, $\widehat{X}$ is the set of maximal vertices of the daisy cube $Q_{n}(X)$ [1], and $0^{n}$ is the minimum vertex of the daisy cube $Q_{n}(X)$. It is well known that daisy cubes are partial cubes. The isometric dimension of a daisy cube $H$ is the minimum dimension of a hypercube that $H$ can be isometrically embedded into.
2.2 Perfect matchings of plane bipartite graphs
A perfect matching of a graph $G$ is a set $M\subseteq E(G)$ of vertex disjoint edges such that every vertex of $G$ is incident to an edge from $M$. A path (or, a cycle) of $G$ is called $M$-alternating if its edges are in and out of $M$ alternately. Moreover, a path (or, a cycle) of $G$ is called $(M_{1},M_{2})$-alternating if its edges are contained in $M_{1}$ and $M_{2}$ alternately for distinct perfect matchings $M_{1}$ and $M_{2}$ of $G$. A path $P$ of $G$ is called a handle if all internal vertices (if exist) of $P$ are degree-2 vertices of $G$, and each end vertex of $P$ has degree at least three in $G$ [5]. It is clear that any handle of $G$ with more than one edge is $M$-alternating for any perfect matching $M$ of $G$. A subgraph $H$ of a graph $G$ is said to be nice if the induced subgraph $G-V(H)$ has a perfect matching.
Let $G$ be a plane graph. A face $s$ of $G$ is a region enclosed by a set of edges of $G$, which forms the periphery of $s$, denoted by $\partial s$. A face $s$ of $G$ is called finite if the periphery of $s$ encloses a finite region, and infinite otherwise. The periphery of the infinite face is called the periphery of $G$. A face $f$ of $G$ is called $M$-resonant if the periphery of $f$ is an $M$-alternating cycle for a perfect matching $M$ of $G$, and $f$ is called $(M_{1},M_{2})$-resonant if its periphery is an $(M_{1},M_{2})$-alternating cycle for two different perfect matching $M_{1}$ and $M_{2}$ of $G$. The inner dual of $G$, denoted by $G^{*}$, is a graph whose vertices are finite faces of $G$ such that two vertices are adjacent in $G^{*}$ if two corresponding finite faces of $G$ have an edge in common. Vertices on the periphery of $G$ are called exterior vertices, and the remaining vertices are interior vertices of $G$. A plane graph whose vertices are all exterior vertices is called an outerplane graph.
Let $G$ be a plane bipartite graph with a perfect matching. Let $S$ be a set of finite faces of $G$. Then $S$ is called an alternating set of $G$ if there exists a perfect matching $M$ of $G$ such that each face in $S$ is $M$-resonant. The Fries number of $G$ is the maximum cardinality of an alternating set of $G$. In addition, if all faces in an alternating set $S$ are pairwise vertex disjoint, then $S$ is called a resonant set of $G$. Moreover, a resonant set of $G$ is called maximal if it is not contained in any other resonant set of $G$. For simplicity, we use $G-S$ to represent the induced subgraph of $G$ obtained by removing all vertices of finite faces from $S$. It is easy to see that if $S$ is an alternating set of $G$, then the induced subgraph $G-S$ is either empty or has a perfect matching. Furthermore, a resonant set $S$ of $G$ is called a canonical resonant set if $G-S$ is empty or has a unique perfect matching.
An edge of a graph with a perfect matching is called allowed if it is contained in a perfect matching of the graph, and forbidden otherwise. A graph is said to be elementary if all its allowed edges form a connected subgraph. For a graph with a perfect matching, each component of its subgraph obtained by removing all forbidden edges is elementary. It was shown that a bipartite graph $G$ is elementary if and only if it is connected and each edge of $G$ is allowed [13]. In addition, a connected plane bipartite graph with more than two vertices is elementary if and only if the periphery of each face (including the infinite face) is an alternating cycle with respect to some perfect matching of the graph [28]. A face $s$ of a plane graph $G$ is said to be a forcing face of $G$ if the induced subgraph of $G$ obtained by deleting all vertices of $s$ is either empty or has a unique perfect matching. By [5], any 2-connected plane bipartite graph with a forcing face is elementary. The concept of a plane weakly elementary bipartite graph was first introduced in [28] for connected graphs. It is more practical to extend the concept to include non-connected graphs too (for example, see [29]). A plane bipartite graph (not necessarily connected) with a perfect matching is called weakly elementary if deleting all forbidden edges does not produce any new finite face. By definition, any plane elementary bipartite graph is also weakly elementary. Note that all allowed edges of a plane weakly elementary bipartite graph $G$ form elementary components of $G$, and each elementary component of $G$ different from $K_{2}$ is called nontrivial.
A benzenoid system is a 2-connected subgraph of the hexagonal lattice such that every finite face is a hexagon. It is well known that every benzenoid system with a perfect matching is weakly elementary [22].
We now present the concept of a reducible face decomposition of a plane elementary bipartite graph with $n$ finite faces [28], which is built on a bipartite ear decomposition of an elementary bipartite graph [13]. A bipartite ear decomposition is defined as follows. Start from an edge $e$, join its two end vertices by a path $P_{1}$ of odd length to obtain an even cycle $G_{1}$, proceed inductively to build a sequence of plane bipartite graphs $G_{i}$ for $2\leq i\leq n$ where $G_{n}=G$ as follows. If $G_{i-1}=e+P_{1}+\ldots+P_{i-1}$ has already been constructed, then $G_{i}=G_{i-1}+P_{i}$ can be obtained by adding the $i$th path $P_{i}$ of odd length such that $P_{i}$ has exactly two end vertices in common with $G_{i-1}$, and $P_{i}$ is called the $i$th ear of $G$. A bipartite ear decomposition is called a reducible face decomposition if $G_{1}$ is the periphery of a finite face $s_{1}$ of $G$, and the $i$th ear $P_{i}$ lies in the exterior of $G_{i-1}$ such that $P_{i}$ and a part of the periphery of $G_{i-1}$ form the periphery of a finite face $s_{i}$ of $G$ for all $i\in\{2,\ldots,n\}$. For such a decomposition, we use notation $\textrm{RFD}(G_{1},G_{2},\ldots,G_{n})$, where $G_{n}=G$. It was shown [28] that a plane bipartite graph different from $K_{2}$ is elementary if and only if it has a reducible face decomposition starting from any finite face of the graph.
Assume that $G$ is a plane bipartite graph whose vertices are properly colored black and white such that adjacent vertices receive different colors. Let $M$ be a perfect matching of $G$. An $M$-alternating cycle $C$ of $G$ is proper $M$-alternating (respectively, improper $M$-alternating) if every edge of $C$ belonging to $M$ goes from white to black vertex (respectively, from black to white vertex) along the clockwise orientation of $C$. A plane bipartite graph $G$ with a perfect matching has a unique perfect matching $M_{\hat{0}}$ (respectively, $M_{\hat{1}}$) such that $G$ has no proper $M_{\hat{0}}$-alternating cycles (respectively, no improper $M_{\hat{1}}$-alternating cycles) [27]. If $G$ is a plane (weakly) elementary bipartite graph, then there is a finite distributive lattice structure on the set $\mathcal{M}(G)$ of all perfect matchings of $G$ [14], and the $height(\mathcal{M}(G))$ is the distance between $M_{\hat{0}}$ and $M_{\hat{1}}$ in the resonance graph $R(G)$ [25].
2.3 Resonance graphs
The resonance graph (also called $Z$-transformation graph) $R(G)$ of a plane bipartite graph $G$ is the graph whose vertices are the perfect matchings of $G$, and two perfect matchings $M_{1},M_{2}$ are adjacent if their symmetric difference $M_{1}\oplus M_{2}$ forms the edge set of exactly one cycle that is the periphery of a finite face $s$ of $G$ [29], and the edge $M_{1}M_{2}$ is said to have the face-label $s$. By definitions, we can see that if $G$ is a plane weakly elementary bipartite graph with elementary components $G_{1},G_{2},\ldots,G_{t}$, then its resonance graph $R(G)$ is the Cartesian product $\Box_{i=1}^{t}R(G_{i})$, see a detail explanation provided in Section 2.3 [1]. For a plane bipartite graph $G$ with a perfect matching, its resonance graph $R(G)$ is connected if and only if $G$ is weakly elementary [29].
In [1], we introduced a new concept called peripherally 2-colorable graphs to characterize when resonance graphs are daisy cubes. Let $G$ be a plane elementary bipartite graph different from $K_{2}$. Then $G$ is called peripherally 2-colorable if every vertex of $G$ has degree 2 or 3, vertices with degree 3 (if exist) are all exterior vertices of $G$, and $G$ can be properly colored black and white so that two vertices with the same color are nonadjacent, and vertices with degree 3 (if exist) are alternatively black and white along the clockwise orientation of the periphery of $G$ (see Figure 1).
Figure 1: A peripherally 2-colorable graph.
In [1], we showed that the resonance graph $R(G)$ of a plane elementary bipartite graph $G$ different from $K_{2}$ is a daisy cube if and only if $G$ is peripherally 2-colorable. Furthermore, the resonance graph $R(G)$ of a plane bipartite graph $G$ is a daisy cube if and only if $G$ is weakly elementary and every elementary component of $G$ different from $K_{2}$ is peripherally 2-colorable.
3 Hypercubes of $R(G)$ and resonant sets of $G$
We start from a relation between resonant sets of a plane weakly elementary bipartite graph and hypercubes of its resonance graph, which generalizes the main result (Theorem 2) in [15] from a benzenoid graph with a perfect matching to any plane weakly elementary bipartite graph.
Lemma 3.1
Let $G$ be a plane weakly elementary bipartite graph different from $K_{2}$. Let $\mathcal{H}(R(G))$ be the set of nontrivial hypercubes of $R(G)$. Let $\mathcal{RS}(G)$ be the set of nonempty resonant sets of $G$. Let $k$ be a positive integer. Define $f:\mathcal{H}(R(G))\rightarrow\mathcal{RS}(G)$ such that if $Q$ is a $k$-dimensional hypercube of $R(G)$, then $f(Q)=S_{Q}$ is a cardinality $k$ resonant set of $G$ with the property that each finite face in $S_{Q}$ is the face-label of a $\Theta$-class of $Q$. Then $f$ is well-defined and surjective. Moreover, $f^{-1}(S)$ is a unique $k$-dimensional hypercube of $R(G)$ only when the cardinality $k$ resonant set $S$ of $G$ is canonical.
Proof.  It is well known that any two edges on a shortest path cannot be in the same $\Theta$-class of a graph. Moreover, for a plane weakly elementary bipartite graph $G$, antipodal edges of a 4-cycle in $R(G)$ have the same face-label, and face-labels of two incident edges of a 4-cycle in $R(G)$ are vertex disjoint finite faces of $G$ [26]. It follows that edges incident to a common vertex of a hypercube in a graph are pairwise contained in different $\Theta$-classes of the graph, and the face-labels of edges incident to a common vertex of a hypercube in $R(G)$ are pairwise vertex disjoint finite faces of $G$.
Let $k$ be a positive integer and $Q$ be a $k$-dimensional hypercube of $R(G)$. Then the edge set of $Q$ consists of $k$ $\Theta$-classes such that each $\Theta$-class has a face-label, and the set of face-labels for the $\Theta$-classes of $Q$ corresponds to a unique set $S_{Q}$ of $k$ vertex disjoint finite faces of $G$. Let $S_{Q}=\{s_{1},s_{2},\ldots,s_{k}\}$. We show that each face in $S_{Q}$ is $M$-resonant for any perfect matching $M$ of $G$ that is a vertex of the hypercube $Q$ of $R(G)$. Starting from $M$, we can construct a shortest path $P$ of length $k$ and contained in $Q$. Then $P$ is a shortest path of $R(G)$, and so the face-labels of edges on $P$ are pairwise distinct since any two edges on $P$ are contained in different $\Theta$-class of $R(G)$. It follows that $S_{Q}$ is the set of face-labels of $k$ edges on $P$. Then there is a permutation $\pi$ on $\{1,2,\ldots,k\}$ such that $M\oplus E(s_{\pi(1)})\oplus E(s_{\pi(2)})\oplus\cdots\oplus E(s_{\pi(k)})$ is the other end vertex of $P$. Recall that all finite faces in $S_{Q}$ are pairwise vertex disjoint, and so $M$-resonant. Therefore, the set of finite faces of $G$ corresponding to the set of face-labels of $\Theta$-classes of $Q$ is a unique cardinality $k$ resonant set $S_{Q}$ of $G$. So, $f$ is well-defined.
Let $S$ be a cardinality $k$ resonant set of $G$ for some positive integer $k$. Then $S$ has $2^{k}$ perfect matchings because the periphery of each finite face in $S$ is a cycle with two perfect matchings. Let $S=\{s_{1},s_{2},\ldots,s_{k}\}$. Let $\mathcal{M}(S)$ be the set of all perfect matchings of $S$. Then $|\mathcal{M}(S)|=2^{k}$ and any $\widehat{m}\in\mathcal{M}(S)$ has the property that all finite faces of $S$ are $\widehat{m}$-resonant. Assume that the vertices of $G$ are properly colored black and white such that adjacent vertices receive different colors. Then for each $\widehat{m}\in\mathcal{M}(S)$, we can assign $\widehat{m}$ a binary string $l_{k}[\widehat{m}]$ of length $k$ as $b_{1}b_{2}\ldots b_{k}$ such that $b_{i}=0$ if $\partial s_{i}$ is proper $\widehat{m}$-alternating, and $b_{i}=1$ otherwise. Then $\{l_{k}[\widehat{m}]\mid\widehat{m}\in\mathcal{M}(S)\}$ induces a $k$-dimensional hypercube $Q_{|S|}$ such that any edge has two end vertices differing in exactly one position $i$ and corresponding to the finite face $s_{i}\in S$, for some $1\leq i\leq k$.
Let $\mathcal{M}(G-S)$ be the set of all perfect matchings of $G-S$. If $G-S$ is empty, then we assume that $\emptyset$ is a perfect matching of $G-S$. Then each perfect matching $m$ in $\mathcal{M}(G-S)$ can be extended to a set of $2^{k}$ perfect matchings of $G$ whose restriction on $S$ is the set of $2^{k}$ perfect matchings $\mathcal{M}(S)$, which induces a $k$-dimensional hypercube $Q^{m}_{|S|}$ of $R(G)$ such that the set of face-labels of $\Theta$-classes of $Q^{m}_{|S|}$ is $S$. Hence, for each cardinality $k$ resonant set $S$ of $G$, there is a set $\mathcal{H}_{S}=\{Q^{m}_{|S|}\mid m\in\mathcal{M}(G-S)\}$ of $k$-dimensional hypercubes of $R(G)$ with the property that for each hypercube $Q^{m}_{|S|}$ in $\mathcal{H}_{S}$, $S$ is the set of face-labels of $\Theta$-classes of $Q^{m}_{|S|}$. It follows that $f^{-1}(S)=\mathcal{H}_{S}$, and so $f$ is surjective.
Note that $|\mathcal{H}_{S}|=|\mathcal{M}(G-S)|$. Then $|f^{-1}(S)|=1$ only when $|\mathcal{M}(G-S)|=1$, that is, $G-S$ is either empty or has a unique perfect matching. Hence, $|f^{-1}(S)|=1$ only when $S$ is a canonical resonant set of $G$. $\Box$
Recall [1] that for a plane elementary bipartite graph $G$ different from $K_{2}$, $R(G)$ is a daisy cube if and only if $G$ is peripherally 2-colorable. By definition, if $G$ is peripherally 2-colorable, then the induced subgraph of $G$ obtained by removing vertices on the periphery of $G$ is either empty or has a unique perfect matching, and so the infinite face of $G$ is forcing. We next consider plane elementary bipartite graphs whose infinite face is forcing.
Lemma 3.2
Let $G$ be a plane elementary bipartite graph different from $K_{2}$. Then the infinite face of $G$ is forcing if and only if any two vertex disjoint cycles whose union forms a nice subgraph of $G$ have disjoint interiors.
Proof.  Sufficiency. Suppose that the infinite face of $G$ is not forcing. Then the resulted subgraph $G^{\prime}$ obtained by deleting all vertices on the periphery cycle $\partial G$ has at least two perfect matchings $m_{1}$ and $m_{2}$. It is well known that the edges in the symmetric difference of $m_{1}$ and $m_{2}$ form a set $S$ of vertex disjoint $(m_{1},m_{2})$-alternating cycles contained in $G^{\prime}$. Let $M_{1}$ and $M_{2}$ be perfect matchings of $G$ extended from $m_{1}$ and $m_{2}$ respectively such that $\partial G$ is $(M_{1},M_{2})$-alternating. Then $S$ is also a set of vertex disjoint $(M_{1},M_{2})$-alternating cycles. Let $C$ be a cycle from $S$. Then the union of $C$ and $\partial G$ forms a nice subgraph of $G$, since both $C$ and $\partial G$ are $(M_{1},M_{2})$-alternating. But the interior of $C$ is contained in the interior of $\partial G$, which is a contradiction.
Necessity. Suppose that there are two vertex disjoint cycles $C_{1}$ and $C_{2}$ whose union forms a nice subgraph of $G$ while their interiors are not disjoint. Then $G$ has a perfect matching $M$ such that both $C_{1}$ and $C_{2}$ are $M$-alternating. Without loss of generality, we can assume that $C_{1}$ is contained in the interior of $C_{2}$. By Corollary 3.4 in [28], there exists a finite face $s_{1}$ of $G$ contained in the interior of $C_{1}$ such that $s_{1}$ is $M$-resonant and vertex disjoint from $C_{2}$. For $i=1,2$, let $I[C_{i}]$ denote the induced plane subgraph of $G$ generated by all vertices either on $C_{i}$ or contained in the interior of $C_{i}$. Note that the restriction of $M$ on $I[C_{2}]$ is a perfect matching $m$ of $I[C_{2}]$ such that both $C_{1}$ and $C_{2}$ are $m$-alternating. It follows that the infinite face of $I[C_{2}]$ is not forcing. By Theorem 2.9 in [28], each $I[C_{i}]$ for $i=1,2$ is elementary. Assume that $I[C_{1}]$ contains $n_{1}$ finite faces and $I[C_{2}]$ contains $n_{2}$ finite faces. Then $n_{1}<n_{2}$, and all finite faces of $I[C_{1}]$ are also finite faces of $I[C_{2}]$. We can construct a RDF$(G_{1},\ldots,G_{n_{1}},G_{n_{1}+1},\ldots,G_{n_{2}},G_{n_{2}+1},\ldots,G_{n})$ of $G=G_{n}$ starting from $s_{1}$ with a sequence of associated finite faces $s_{1},\ldots,s_{n_{1}},s_{n_{1}+1},\ldots,s_{n_{2}},s_{n_{2}+1},\ldots,s_{n}$ such that $I[C_{1}]$ has a RDF$(G_{1},\ldots,G_{n_{1}})$ with a sequence of associated finite faces $s_{1},\ldots,s_{n_{1}}$, and $I[C_{2}]$ has a RDF$(G_{1},\ldots G_{n_{1}},G_{n_{1}+1},\ldots,G_{n_{2}})$ with a sequence of associated finite faces $s_{1},\ldots,s_{n_{1}},s_{n_{1}+1},\ldots,s_{n_{2}}$. Then by the proof of Theorem 2 in [4], $height(\mathcal{M}(G_{1}))=1$, and for $2\leq i\leq n$, $height(\mathcal{M}(G_{i}))\geq height(\mathcal{M}(G_{i-1}))+1$. Moreover, for each $2\leq i\leq n$, $height(\mathcal{M}(G_{i}))\geq i$ and the equality holds if and only if the infinite face of $G_{i}$ is forcing. Recall that the infinite face of $G_{n_{2}}=I[C_{2}]$ is not forcing. It follows that $height(\mathcal{M}(G))>n$, and so the infinite face of $G$ is not forcing. $\Box$
A coronene is the benzenoid system shown in Figure 2. It is known [23] that an elementary benzenoid system $G$ has no coronenes as nice subgraphs if and only if any two vertex disjoint cycles whose union forms a nice subgraph of $G$ have disjoint interiors. The following corollary follows immediately by Lemma 3.2.
Figure 2: A coronene.
Corollary 3.3
Let $G$ be an elementary benzenoid system. Then the infinite face of $G$ is forcing if and only if $G$ has no coronenes as nice subgraphs.
Next, we present a bijection between the set of maximal hypercubes of $R(G)$ and the set of maximal resonant sets of a plane elementary bipartite graph $G$ whose infinite face is forcing, which extends Lemma 4.2 and Theorem 4.3 in [18] from an elementary benzenoid system without coronenes as nice subgraphs to a plane elementary bipartite graph whose infinite face is forcing.
Theorem 3.4
Let $G$ be a plane elementary bipartite graph whose infinite face is forcing. Then the following properties hold true.
  • $(i)$
    Let $S$ be a resonant set of $G$. Then $S$ is maximal if and only if $S$ is canonical.
  • $(ii)$
    Let $\mathcal{H}(R(G))$ be the set of nontrivial hypercubes of $R(G)$, $\mathcal{RS}(G)$ be the set of nonempty resonant sets of $G$, and $f:\mathcal{H}(R(G))\rightarrow\mathcal{RS}(G)$ be the mapping defined in Lemma 3.1. Then $f$ induces a bijection between the set of maximal hypercubes of $R(G)$ and the set of maximal resonant sets of $G$ which maps a maximal hypercube $Q$ of $R(G)$ to a maximal resonant set $S$ of $G$ such that $S$ is the set of face-labels of $\Theta$-classes of $Q$.
Proof.  $(i)$ Sufficiency. Let $S$ be a canonical resonant set of $G$. Then $G-S$ is either empty or has a unique perfect matching. Hence, if $M$ is a perfect matching of $G$ such that each finite face in $S$ is $M$-resonant, then $G$ has no other $M$-resonant faces except the faces in $S$. It implies that $S$ cannot be contained in any other resonant set of $G$, and so $S$ is maximal.
Necessity. Let $S$ be a maximal resonant set of $G$. Suppose that $S$ is not canonical. Then $G-S$ has at least two different perfect matchings $m_{1}$ and $m_{2}$. Since the symmetric difference of any two perfect matchings is a set of edges of vertex disjoint cycles, it follows that $G-S$ contains at least one $(m_{1},m_{2})$-alternating cycle $C$. Let $M_{1}$ and $M_{2}$ be two perfect matchings of $G$ that are extended from $m_{1}$ and $m_{2}$, respectively, such that all finite faces contained in $S$ are $(M_{1},M_{2})$-resonant. Then the $(m_{1},m_{2})$-alternating cycle $C$ contained in $G-S$ is also an $(M_{1},M_{2})$-alternating cycle. Let $I[C]$ denote the induced plane subgraph of $G$ generated by all vertices either on $C$ or contained in the interior of $C$. Then by Corollary 3.4 in [28], there exists a finite face $s_{0}$ of $G$ that is contained in $I[C]$, and $s_{0}$ is $M_{1}$-resonant. If $S$ and $I[C]$ are vertex disjoint, then $S\cup\{s_{0}\}$ is a resonant set. This is a contradiction to the assumption that $S$ is maximal. If $S$ and $I[C]$ are not vertex disjoint, then there exists a finite face $s\in S$ contained in $I[C]$. Note that $C$ cannot be the periphery of any finite face in $S$ since $C$ is contained in $G-S$. It follows that the periphery of $s\in S$ and $C$ are two vertex disjoint cycles which form a nice subgraph of $G$, but the interior of $s$ is contained in the interior of $C$. Then by Lemma 3.2, the infinite face of $G$ is not forcing. This is a contradiction.
$(ii)$ Let $f:\mathcal{H}(R(G))\rightarrow\mathcal{RS}(G)$ be a mapping defined in Lemma 3.1, where $\mathcal{H}(R(G))$ is the set of nontrivial hypercubes of $R(G)$ and $\mathcal{RS}(G)$ is the set of nonempty resonant sets of $G$. Let $S$ be a maximal resonant set of $G$ of cardinality $k$, where $k$ is a positive integer. By $(i)$, $S$ is canonical. Lemma 3.1 then assures that $f^{-1}(S)$ is a unique $k$-dimensional hypercube $Q$ from $\mathcal{H}(R(G))$ such that $S$ is the set of face-labels of $\Theta$-classes of $f^{-1}(S)=Q$. If $Q$ is not a maximal hypercube of $R(G)$, then there exists a $k^{\prime}$-dimensional hypercube $Q^{\prime}$ of $R(G)$ such that $k^{\prime}>k$ and $Q$ is contained in $Q^{\prime}$. It follows that $f(Q^{\prime})=S^{\prime}$ is a cardinality $k^{\prime}$ resonant set of $G$ such that $S\subseteq S^{\prime}$. This is a contradiction since $S$ is a maximal resonant set. We have shown that if $S$ is a cardinality $k$ maximal resonant set of $G$, then $f^{-1}(S)$ is a $k$-dimensional maximal hypercube of $R(G)$ such that $S$ is the set of face-labels of $\Theta$-classes of $f^{-1}(S)$.
Next, we prove that if $Q$ is a $k$-dimensional maximal hypercube of $R(G)$, then $f(Q)=S$ is a cardinality $k$ maximal resonant set of $G$ consisting of the face-labels of the $\Theta$-classes of $Q$. Let $S=\{s_{1},s_{2},\ldots,s_{k}\}$. Let $\mathcal{M}(G-S)$ be the set of all perfect matchings of $G-S$. By the proof of Lemma 3.1, we can let $\mathcal{H}_{S}=\{Q^{m}_{|S|}\mid m\in\mathcal{M}(G-S)\}$ be the set of $k$-dimensional hypercubes in $R(G)$ with the property that for each hypercube $Q^{m}_{|S|}$ in $\mathcal{H}_{S}$, $S$ is the set of face-labels of $\Theta$-classes of $Q^{m}_{|S|}$. Then $Q=Q^{m_{1}}_{|S|}$ for some $m_{1}\in\mathcal{M}(G-S)$. So, any vertex of $Q$ is a perfect matching $M$ of $G$ such that $M\cap E(G-S)=m_{1}$ and all finite faces in $S$ are $M$-resonant.
Suppose that $S$ is not maximal. Then $S$ is not canonical by $(i)$. Then $G-S$ has at least one other perfect matching $m_{2}$ that is different from $m_{1}$. Let $M_{1}$ be a perfect matching of $G$ such that $M_{1}\cap E(G-S)=m_{1}$ and all finite faces in $S$ are $M_{1}$-resonant. Let $M_{2}$ be a perfect matching of $G$ such that $M_{2}\cap E(G-S)=m_{2}$ and the restriction of $M_{2}$ on $S$ is identical to the restriction of $M_{1}$ on $S$. Then $M_{1}\oplus M_{2}=m_{1}\oplus m_{2}$ is a set of $(M_{1},M_{2})$-alternating cycles contained in $G-S$. Let $C$ be an $(M_{1},M_{2})$-alternating cycle that lies in $G-S$ and let $I[C]$ be the induced plane subgraph of $G$ generated by all vertices on $C$ or in the interior of $C$. By Corollary 3.4 [28], there is a finite face $s_{k+1}$ of $G$ contained in $I[C]$ that is also $M_{1}$-resonant. It is clear that $s_{k+1}$ is also $m_{1}$-resonant. Recall that the infinite face of $G$ is forcing. If $C$ is vertex disjoint with the periphery of $s_{k+1}$, then the union $C$ and $\partial s_{k+1}$ forms a nice subgraph of $G$, while interior of $s_{k+1}$ is contained in the interior of $C$. This is a contradiction to the conclusion of Lemma 3.2. Therefore, $C$ must have nonempty intersection with $s_{k+1}$. Moreover, $s_{k+1}\notin S$ since $C$ belongs to $G-S$. Let $S^{\prime}=S\cup\{s_{k+1}\}$. Recall that $m_{1}\in\mathcal{M}(G-S)$ and $s_{k+1}$ is $m_{1}$-resonant. Let $m^{\prime}_{2}=m_{1}\oplus E(s_{k+1})$. Then $m^{\prime}_{2}\in\mathcal{M}(G-S)$. (It is possible that $m^{\prime}_{2}=m_{2}$.)
Let $Q^{\prime}=Q^{m^{\prime}_{2}}_{|S|}\in\mathcal{H}_{S}$. Then any vertex of $Q^{\prime}$ is a perfect matching $M^{\prime}$ of $G$ such that $M^{\prime}\cap E(G-S)=m^{\prime}_{2}$ and all finite faces in $S$ are $M^{\prime}$-resonant. Note that $m_{1}\oplus m^{\prime}_{2}=E(s_{k+1})$. Recall that any vertex of $Q$ is a perfect matching $M$ of $G$ such that $M\cap E(G-S)=m_{1}$ and all finite faces in $S$ are $M$-resonant. Hence, for each vertex $M$ of the hypercube $Q$, there is a unique vertex $M^{\prime}$ of the hypercube $Q^{\prime}$ such that $M^{\prime}\cap E(G-S)=m^{\prime}_{2}$ and the restriction of $M^{\prime}$ on $S$ is identical to the restriction of $M$ on $S$. It follows that $M\oplus M^{\prime}=m_{1}\oplus m^{\prime}_{2}=E(s_{k+1})$. By definition, $MM^{\prime}$ is an edge of $R(G)$ with the face-label $s_{k+1}$. Note that the symmetric difference of any two perfect matchings that are vertices in the hypercube $Q$ contains the edges of the periphery of least one face from $S$, and $s_{k+1}$ is vertex disjoint from any finite face from $S$. It follows that $M^{\prime}$ is the unique vertex of $Q^{\prime}$ that is adjacent to the given vertex $M$ of $Q$ in $R(G)$, since two perfect matchings are adjacent in $R(G)$ if their symmetric difference is the periphery of exactly one finite face of $G$. Let $V(Q)$ and $V(Q^{\prime})$ be the vertex set of $Q$ and $Q^{\prime}$ respectively. Then there is a matching between $V(Q)$ and $V(Q^{\prime})$ in $R(G)$ such that each edge of the matching has the face-label $s_{k+1}$. Hence, the induced subgraph of $R(G)$ on $V(Q)\cup V(Q^{\prime})$ is a $(k+1)$-dimensional hypercube of $R(G)$ which contains $Q$. This is a contradiction to the assumption that $Q$ is maximal.
Therefore, $f$ induces a bijection between the set of maximal hypercubes of $R(G)$ and the set of maximal resonant sets of $G$ which maps a maximal hypercube $Q$ of $R(G)$ to a maximal resonant set $S$ of $G$ such that $S$ is the set of face-labels of $\Theta$-classes of $Q$. $\Box$
4 A characterization on when $R(G)$ is a daisy cube
Recall that [1] for a plane elementary bipartite graph $G$ different from $K_{2}$, $R(G)$ is daisy cube if and only if $G$ is peripherally 2-colorable. We first provide a connection between (maximal) resonant sets of a peripherally 2-colorable graph $G$ and (maximal) independent sets of its inner dual $G^{*}$.
Lemma 4.1
Let $G$ be a peripherally 2-colorable graph and $G^{*}$ be the inner dual of $G$. Let $\mathcal{RS}(G)$ be the set of nonempty resonant sets of $G$, and $\mathcal{IS}(G^{*})$ be the set of nonempty independent sets of $G^{*}$. Define $g:\mathcal{RS}(G)\rightarrow\mathcal{IS}(G^{*})$ such that $g(S)=S^{*}$, where $S$ is a resonant set of $G$, and $S^{*}$ is the corresponding independent set of $G^{*}$. Then $g$ is a well-defined bijection which maps a resonant set $S$ of $G$ to an independent set $S^{*}$ of $G^{*}$ where $|S|=|S^{*}|=k$ for some positive integer $k$. Moreover, $S$ is a maximal resonant set of $G$ if and only if $S^{*}$ is a maximal independent set of $G^{*}$.
Proof.  If $S$ is a resonant set of $G$, then by definition, the finite faces of $G$ contained in $S$ are pairwise vertex disjoint. It follows that its corresponding vertex set $S^{*}$ contained in $G^{*}$ is an independent set of $G^{*}$, and so $g$ is well-defined. We first show that $g$ is a bijection. It is trivial that $g$ is injective since different resonant sets of $G$ have different corresponding vertex sets in $G^{*}$. On the other hand, if $S^{*}$ is an independent set of $G^{*}$, then the finite faces in the corresponding set $S$ of $G$ are pairwise edge disjoint. Since $G$ is peripherally 2-colorable, it follows that any two finite faces in $S$ are vertex disjoint because any vertex degree of $G$ is at most 3. By the definition of a peripherally 2-colorable graph, each handle of $G$ is a path on an even number of vertices, which implies that the induced subgraph $G-S$ is either empty or has a perfect matching. Therefore, $S$ is a resonant set of $G$ and so $g$ is surjective.
We then show that $S$ is a maximal resonant set of $G$ if and only if $S^{*}=g(S)$ is a maximal independent set of $G^{*}$. Necessity. Let $S$ be a maximal resonant set of $G$. Suppose that $S^{*}$ is not maximal. Then there exists an independent set $T^{*}$ of the inner dual $G^{*}$ such that $S^{*}\subsetneq T^{*}$. Let $T$ be the set of finite faces of $G$ corresponding to the vertices from $T^{*}$. Similarly to the above paragraph, we can show that $T$ is also a resonant set of $G$, which is a contradiction to the assumption that $S$ is maximal. Sufficiency. Assume that $S^{*}$ is a maximal independent set of the inner dual $G^{*}$. Then the corresponding set $S$ of finite faces in $G$ is a resonant set of $G$. Therefore, the induced subgraph $G-S$ is either empty or has a perfect matching. Suppose that $S$ is not maximal. Then there exists a resonant set $T$ such that $S\subsetneq T$. However, its corresponding set $T^{*}$ of vertices in the inner dual $G^{*}$ is obviously an independent set of $G^{*}$ and $S^{*}\subsetneq T^{*}$, which is a contradiction to the assumption that $S^{*}$ is maximal.
It is clear that if $g(S)=S^{*}$, then $|S|=|S^{*}|=k$ for some positive integer $k$. Therefore, the conclusion follows. $\Box$
Based on Theorem 3.4 and Lemma 4.1, for a peripherally 2-colorable graph $G$, we can establish a bijection between the set of maximal hypercubes of its resonance graph $R(G)$ and the set of maximal independent sets of its inner dual $G^{*}$.
Corollary 4.2
Let $G$ be a peripherally 2-colorable graph and $G^{*}$ be the inner dual of $G$. Moreover, let $f:\mathcal{H}(R(G))\rightarrow\mathcal{RS}(G)$ be a mapping defined in Lemma 3.1 and $g:\mathcal{RS}(G)\rightarrow\mathcal{IS}(G^{*})$ a mapping defined in Lemma 4.1. Define $h:\mathcal{H}(R(G))\rightarrow\mathcal{IS}(G^{*})$ such that $h(Q)=g(f(Q))$ for any $Q\in\mathcal{H}(R(G))$. Then $h$ induces a bijection between the set of maximal hypercubes of $R(G)$ and the set of maximal independent sets of $G^{*}$ such that $h(Q)=S_{Q}^{*}$ if $Q$ is a $k$-dimensional maximal hypercube of $R(G)$, and $S_{Q}^{*}$ is a cardinality $k$ maximal independent set of $G^{*}$ with the property that each vertex of $S_{Q}^{*}$ corresponds to the face-label of a unique $\Theta$-class of $Q$, where $k$ is a positive integer.
Proof.  Recall that the infinite face of any peripherally 2-colorable graph is forcing. Then the conclusion on the bijection is trivial by Theorem 3.4 and Lemma 4.1. $\Box$
Note that Corollary 4.2 can be generalized to plane weakly elementary bipartite graphs whose each elementary component different from $K_{2}$ is peripherally 2-colorable.
We now show that for any graph $H$, we can construct a daisy cube from the set of all independent sets (including the empty set) of $H$.
Lemma 4.3
Let $H$ be a graph with the vertex set $\{v_{1},v_{2},\ldots,v_{n}\}$ and let $\mathcal{IS}_{0}(H)$ be the set of all independent sets of $H$ (including the empty set). For an arbitrary $A\in\mathcal{IS}_{0}(H)$, define $\chi(A)=\chi_{A}(v_{1})\chi_{A}(v_{2})\ldots\chi_{A}(v_{n})$, where $\chi_{A}(v_{j})=1$ if $v_{j}\in A$ and $\chi_{A}(v_{j})=0$ otherwise. Let $D_{\mathcal{I}}(H)=\langle\{\chi(A)\in\mathcal{B}^{n}\mid A\in\mathcal{IS}_{0}(H)\}\rangle$ be an induced subgraph of $Q_{n}$. Then $D_{\mathcal{I}}(H)$ is a daisy cube. Moreover, if =$\mathcal{I}=\{I_{1},I_{2},\ldots,I_{t}\}$ is the set of maximal independent sets of $H$, then $X_{\mathcal{I}}=\{\chi(I_{1}),\chi(I_{2}),\ldots,\chi(I_{t})\}$ is the set of maximal vertices of $D_{\mathcal{I}}(H)$, and $D_{\mathcal{I}}(H)=Q_{n}(X_{\mathcal{I}})=\langle\{u\in\mathcal{B}^{n}\mid u\leq\chi(I_{i})\textrm{ for some }\chi(I_{i})\in X_{\mathcal{I}}\}\rangle$.
Proof.  Note that the vertex set $V(D_{\mathcal{I}}(H))=\{\chi(A)\in\mathcal{B}^{n}\mid A\in\mathcal{IS}_{0}(H)\}$ is contained in the poset $(\mathcal{B}^{n},\leq)$ such that $\chi(A_{1})\leq\chi(A_{2})$ in $\mathcal{B}^{n}$ if $A_{1}\subseteq A_{2}$. By definition, for the poset $(V(D_{\mathcal{I}}(H)),\leq)$, we can see that $\chi(\emptyset)=0^{n}$ is the minimum vertex of $D_{\mathcal{I}}(H)$. Moreover, if =$\mathcal{I}=\{I_{1},I_{2},\ldots,I_{t}\}$ is the set of maximal independent sets of $H$, then $X_{\mathcal{I}}=\{\chi(I_{1}),\chi(I_{2}),\ldots,\chi(I_{t})\}$ is the set of maximal elements of the poset $(V(D_{\mathcal{I}}(H)),\leq)$. Hence, $X_{\mathcal{I}}$ is the set of maximal vertices of $D_{\mathcal{I}}(H)$.
For $1\leq i\leq t$, let $\mathcal{P}(I_{i})$ be the set of subsets (including the empty set) of the maximal independent set $I_{i}$ of $H$. Then $\mathcal{IS}_{0}(H)=\cup_{i=1}^{t}\mathcal{P}(I_{i})$. Obviously, $V(D_{\mathcal{I}}(H))=\{\chi(A)\mid A\in\cup_{i=1}^{t}\mathcal{P}(I_{i})\}=\cup_{i=1}^{t}\{\chi(A)\mid A\in\mathcal{P}(I_{i})\}$. Since $D_{\mathcal{I}}(H)$ is an induced subgraph of $Q_{n}$, it follows that for each $1\leq i\leq t$, the induced subgraph $\langle\{\chi(A)\mid A\in\mathcal{P}(I_{i})\}\rangle=\langle\{u\in\mathcal{B}^{n}\mid u\leq\chi(I_{i})\}\rangle$ is a hypercube of $D_{\mathcal{I}}(H)$ which contains $\chi(\emptyset)=0^{n}$ since $\mathcal{P}(I_{i})$ contains the empty set.
Note that $V(D_{\mathcal{I}}(H))=\cup_{i=1}^{t}\{\chi(A)\mid A\in\mathcal{P}(I_{i})\}=\cup_{i=1}^{t}\{u\in\mathcal{B}^{n}\mid u\leq\chi(I_{i})\}=\{u\in\mathcal{B}^{n}\mid u\leq\chi(I_{i})\,\text{ for some }\,i\in\{1,2,\ldots,t\}\}$. Recall that $D_{\mathcal{I}}(H)$ is an induced subgraph of $Q_{n}$ and $X_{\mathcal{I}}=\{\chi(I_{1}),\chi(I_{2}),\ldots,\chi(I_{t})\}$ is the set of maximal vertices of $D_{\mathcal{I}}(H)$. Then by the definition of a daisy cube, $D_{\mathcal{I}}(H)$ is a daisy cube isomorphic to $Q_{n}(X_{\mathcal{I}})=\langle\{u\in\mathcal{B}^{n}\mid u\leq\chi(I_{i})\,\text{ for some }\,\chi(I_{i})\in X_{\mathcal{I}}\}\rangle$. $\Box$
We then present a characterization on when the resonance graph of a plane elementary bipartite graph $G$ is a daisy cube in terms of the daisy cube generated from the set of all independent sets of its inner dual $G^{*}$.
Lemma 4.4
Let $G$ be a plane elementary bipartite graph and $G^{*}$ be the inner dual of $G$. Then $R(G)$ is a daisy cube if and only if $R(G)$ is isomorphic to $D_{\mathcal{I}}(G^{*})$, where $D_{\mathcal{I}}(G^{*})$ is the graph defined in Lemma 4.3.
Proof.  Let $n$ be the number of finite faces of $G$. If $n=0$, then $G$ is $K_{2}$, and the resonance graph $R(G)$ is the one-vertex graph. On the other hand, the empty set is the only independent set of $G^{*}$, which means that $D_{\mathcal{I}}(G^{*})$ is also the one-vertex graph. In the following we assume that $n>0$.
Sufficiency follows by Lemma 4.3. It remains to show necessity. Suppose that $R(G)$ is a daisy cube. Then by Corollary 3.6 in [1], $G$ is a peripherally 2-colorable graph and the isometric dimension of $R(G)$ is $n$. Hence, we can assume that the vertices of $R(G)$ are binary strings of length $n$. By Lemma 3.4 in [1], there is a 1–1 correspondence between the set of finite faces of $G$ and the set of $\Theta$-classes of $R(G)$ such that each $\Theta$-class of $R(G)$ has a unique face-label which is a finite face of $G$. Moreover, each $\Theta$-class of $R(G)$ has exactly one edge incident to $0^{n}$. Note that $R(G)$ is an isometric subgraph of the $n$-dimensional hypercube $Q_{n}$. It follows that for each $j\in\{1,\ldots,n\}$, the position $j$ in the binary string of vertices of $R(G)$ corresponds to a unique finite face of $G$, which we denote by $s_{j}$.
Let $X=\{x_{1},x_{2},\ldots,x_{t}\}\subseteq\mathcal{B}^{n}$ be the set of maximal vertices of $R(G)$. Then $R(G)$ is isomorphic to $Q_{n}(X)=\langle\{u\in\mathcal{B}^{n}\mid u\leq x_{i}\ \text{ for some }\ x_{i}\in X\}\rangle$, and so any maximal hypercube of $R(G)$ is $\langle\{u\in\mathcal{B}^{n}\mid u\leq x_{i}\}\rangle$ for some $x_{i}\in X$. Moreover, for each maximal vertex $x_{i}$ of $R(G)$, the $j$th position (where $1\leq j\leq n$) of $x_{i}$ is 1 if $s_{j}$ is a face-label of an edge in the hypercube $\langle\{u\in\mathcal{B}^{n}\mid u\leq x_{i}\}\rangle$, and 00 otherwise.
By Corollary 4.2, there is a bijection between the set of maximal hypercubes of $R(G)$ and the set of maximal independent sets of $G^{*}$, which maps a maximal hypercube of $R(G)$ to a maximal independent set of $G^{*}$ that corresponds to the set of the face-labels of the $\Theta$-classes of the maximal hypercube. Then we can let $\{I^{*}_{1},I^{*}_{2},\ldots,I^{*}_{t}\}$ be the set of maximal independent sets of $G^{*}$ such that $\{Q_{|I^{*}_{1}|},Q_{|I^{*}_{2}|},\ldots,Q_{|I^{*}_{t}|}\}$ is the set of maximal hypercubes of $R(G)$, where $Q_{|I^{*}_{i}|}=\langle\{u\in\mathcal{B}^{n}\mid u\leq x_{i}\}\rangle$, and the maximal independent set $I^{*}_{i}$ of $G^{*}$ corresponds to the set of face-labels of the $\Theta$-classes of the maximal hypercube $Q_{|I^{*}_{i}|}$ in $R(G)$.
Let $\{s^{*}_{1},s^{*}_{2},\ldots,s^{*}_{n}\}$ be the vertex set of $G^{*}$, where $s^{*}_{j}$ corresponds to the finite face $s_{j}$ of $G$ for $j\in\{1,2,\ldots,n\}$. For each $i\in\{1,2,\ldots,t\}$, define $\chi(I_{i}^{*})\in\mathcal{B}^{n}$ such that the $j$th position (where $1\leq j\leq n$) of $\chi(I_{i}^{*})$ is $1$ if $s^{*}_{j}\in I^{*}_{i}$, and 00 otherwise. It follows that $\chi(I_{i}^{*})=x_{i}$ for $i\in\{1,2,\ldots,t\}$. Hence, $\{\chi(I^{*}_{1}),\chi(I^{*}_{2}),\ldots,\chi(I^{*}_{t})\}=\{x_{1},x_{2},\ldots,x_{t}\}$ is the set of maximal vertices of $R(G)$.
Recall that $\{I^{*}_{1},I^{*}_{2},\ldots,I^{*}_{t}\}$ is the set of maximal independent sets of $G^{*}$. By Lemma 4.3, $\{\chi(I^{*}_{1}),\chi(I^{*}_{2}),\ldots,\chi(I^{*}_{t})\}$ is the set of maximal vertices of $D_{\mathcal{I}}(G^{*})$. Consequently, daisy cubes $D_{\mathcal{I}}(G^{*})$ and $R(G)$ have the same set of maximal vertices, which implies that they are isomorphic. $\Box$
See Figure 3 for the illustration of Corollary 4.2 and Lemma 4.4.
Figure 3: A peripherally 2-colorable graph $G$ whose inner dual $G^{*}$ has three maximal independent sets which correspond to the three maximal vertices of $R(G)$.
We now present a generalized version of Lemma 4.4.
Theorem 4.5
Let $G$ be a plane weakly elementary bipartite graph, $G_{a}$ be the subgraph of $G$ obtained by removing all forbidden edges from $G$, and $G_{a}^{*}$ be the inner dual of $G_{a}$. Then $R(G)$ is a daisy cube if and only if $R(G)$ is isomorphic to $D_{\mathcal{I}}(G_{a}^{*})$, where $D_{\mathcal{I}}(G_{a}^{*})$ is the graph defined in Lemma 4.3.
Proof.  Let $G$ be a plane weakly elementary bipartite graph and $G_{a}$ be the subgraph of $G$ obtained by removing all forbidden edges from $G$ (note that all the edges in $G_{a}$ are allowed). Then $G_{a}$ is a disjoint union of elementary components of $G$, which are denoted by $G_{1},G_{2},\ldots,G_{t}$. Let $G^{*}_{i}$ be the inner dual of $G_{i}$ for $1\leq i\leq t$.
It is known [1] that $R(G)=\Box_{i=1}^{t}R(G_{i})$ and $R(G)$ is a daisy cube if and only if $R(G_{i})$ is a daisy cube for all $1\leq i\leq t$. By Lemma 4.4, $R(G_{i})$ is a daisy cube if and only if $R(G_{i})$ is isomorphic to $D_{\mathcal{I}}(G^{*}_{i})$. Therefore, $R(G)$ is a daisy cube if and only if $R(G)$ is isomorphic to $\Box_{i=1}^{t}D_{\mathcal{I}}(G^{*}_{i})$. By Lemma 4.3, $D_{\mathcal{I}}(G^{*}_{i})$ is a daisy cube such that $\chi(I)$ is a maximal vertex of $D_{\mathcal{I}}(G^{*}_{i})$ if and only if $I$ is a maximal independent set of $G^{*}_{i}$. By the proof of Theorem 4.1 in [1], we can see that $\Box_{i=1}^{t}D_{\mathcal{I}}(G^{*}_{i})$ is a daisy cube such that any maximal vertex of $\Box_{i=1}^{t}D_{\mathcal{I}}(G^{*}_{i})$ is of the form $\chi(I_{1})\chi(I_{2})\ldots\chi(I_{t})$ where $I_{i}$ is a maximal independent set of $G^{*}_{i}$ for $1\leq i\leq t$.
Let $G^{*}_{a}$ be the inner dual of $G_{a}$. By Lemma 4.3, $D_{\mathcal{I}}(G^{*}_{a})$ is a daisy cube such that $\chi(I_{a})$ is a maximal vertex of $D_{\mathcal{I}}(G^{*}_{a})$ if and only if $I_{a}$ is a maximal independent set of $G^{*}_{a}$. Note that any maximal independent set of $G^{*}_{a}$ is a disjoint union of maximal independent sets of $G^{*}_{1},G^{*}_{2},\ldots,G^{*}_{t}$. It follows that any maximal vertex of $D_{\mathcal{I}}(G^{*}_{a})$ is $\chi(I_{1}\cup I_{2}\cup\ldots\cup I_{t})$ where $I_{i}$ is a maximal independent set of $G^{*}_{i}$ for $1\leq i\leq t$. By definition, we can see that $\chi(I_{1})\chi(I_{2})\ldots\chi(I_{t})=\chi(I_{1}\cup I_{2}\cup\cdots\cup I_{t})$. Then two daisy cubes $\Box_{i=1}^{t}D_{\mathcal{I}}(G^{*}_{i})$ and $D_{\mathcal{I}}(G^{*}_{a})$ are isomorphic since they have the same set of maximal vertices.
Therefore, $R(G)$ is a daisy cube if and only if $R(G)$ is isomorphic to $D_{\mathcal{I}}(G_{a}^{*})$. $\Box$
The following corollary follows from the above result.
Corollary 4.6
Let $G$ be a plane weakly elementary bipartite graph whose elementary components different from $K_{2}$ are peripherally 2-colorable. Let $G_{a}$ be the subgraph of $G$ obtained by removing all forbidden edges from $G$, and $G_{a}^{*}$ be the inner dual of $G_{a}$. Then there exists a bijection between the set of all perfect matchings of $G$ and the set of all independent sets of $G^{*}_{a}$. In particular, if $G$ is a peripherally 2-colorable graph and $G^{*}$ is the inner dual of $G$, then there exists a bijection between the set of all perfect matchings of $G$ and the set of all independent sets of $G^{*}$.
Proof.  By [1], $R(G)$ is a daisy cube. By Theorem 4.5, $R(G)$ is isomorphic to $D_{\mathcal{I}}(G^{*}_{a})$. Note that the vertex set of $R(G)$ is the set of all perfect matchings of $G$. By Lemma 4.3, there is a bijection between the vertex set of $D_{\mathcal{I}}(G^{*}_{a})$ and the set of all independent sets of $G^{*}_{a}$. Therefore, there exists a bijection between the set of all perfect matchings of $G$ and the set of all independent sets of $G^{*}_{a}$. In particular, if $G$ is a peripherally 2-colorable graph, then $G^{*}_{a}$ is $G^{*}$, and so the conclusion follows. $\Box$
Note that one specific bijection between the set of all perfect matchings of a peripherally 2-colorable graph $G$ and the set of all independent sets of $G^{*}$ was described in [19].
5 Algorithms
In this section, we present two algorithms which produce a proper labelling for the vertex set of the resonance graph of a peripherally 2-colorable graph.
Algorithm 1 can be considered as the first part of algorithm [3] for the case of an outerplane bipartite graphs being “regular”, but needs to be proved that it generates all perfect matchings with the described properties. Recently, Che and Chen [6] gave an equivalent algorithm together with a proof based on a decomposition structure on a peripherally 2-colorable graph.
Input: $RFD(G_{1},G_{2},\ldots,G_{n})$ of $G=G_{n}$ associated with a sequence $s_{1},\ldots,s_{n}$ of finite faces.
Output: A binary coding of length $n$ for all perfect matchings of $G$.
1 $B:=\{00,01,10\}$
2 for $r=2,\ldots,n-1$ do
3       $B^{\prime}:=\emptyset$
4       set $j\in\left\{1,\ldots,r\right\}$ such that $s_{j}$ is adjacent to $s_{r+1}$
5       for each $x\in B$ do
6            $B^{\prime}:=B^{\prime}\cup\{x0\}$
7             if $x_{j}=0$ then
8                  $B^{\prime}:=B^{\prime}\cup\{x1\}$
9             end if
10
11       end for
12       $B:=B^{\prime}$
13 end for
return $B$
Algorithm 1 A binary coding for all perfect matchings of a peripherally 2-colorable graph $G$ by a reducible face decomposition of $G$.
Next, we focus on the other algorithm. For the sake of completeness, we prove the following lemma (note that this result already follows from the discussion in [19]).
Lemma 5.1
Let $G$ be a peripherally 2-colorable graph and $G^{*}$ be the inner dual of $G$. Then $G^{*}$ is a tree.
Proof.  Note that a peripherally 2-colorable graph $G$ is either a 2-connected outerplane bipartite graph, or can be transformed into a 2-connected outerplane bipartite graph $G^{\prime}$ by the proof of Theorem 3.5 in [1] such that the resulted 2-connected outerplane bipartite graph $G^{\prime}$ is also peripherally 2-colorable, the inner dual of $G^{\prime}$ is isomorphic to the inner dual of $G$, and $R(G^{\prime})$ is isomorphic to $R(G)$. It is clear that the inner dual of any 2-connected outerplane bipartite graph is a tree. So, the inner dual of any peripherally 2-colorable graph is a tree. $\Box$
Algorithm 2 is obtained by Lemma 4.4. In order to do that, we need to find all maximal independent sets of a tree. Wilf [21] showed that for a tree of order $n\geq 1$, the largest number of its maximal independent sets is $12^{n/2-1}+$ 1 or $2^{(n-1)/2}$ depending on whether $n$ is even or odd. He also provided a linear time algorithm for the computation of the number of maximal independent sets of any given tree. Chang et al. [7] presented an algorithm that generates all maximal independent sets for trees of order $n$ in lexicographic order with $O(n^{2})$ delay between the output of any two successive independent sets and with $O(n)$ space.
Input: Inner dual $G^{*}$ with vertices $s^{*}_{1},\ldots,s^{*}_{n}$.
Output: A binary coding for all perfect matchings of $G$.
1 find the maximal independent sets $I_{1},\ldots,I_{k}$ of $G^{*}$
2 $B:=\emptyset$
3 for $r=1,\ldots,k$ do
4       for every subset $X$ of $I_{r}$ do
5             $B:=B\cup\{x_{1}x_{2}\ldots x_{n}\}$, where for every $i\in\{1,\ldots,n\}$, $x_{i}=1$ if $s^{*}_{i}\in X$ and $x_{i}=0$ if $s^{*}_{i}\notin X$
6
7       end for
8
9 end for
return $B$
Algorithm 2 A binary coding for all perfect matchings of a peripherally 2-colorable graph $G$ by the maximal independent sets of the inner dual $G^{*}$ of $G$.
Remark 5.2
If $G$ is a plane bipartite graph whose resonance graph $R(G)$ is a daisy cube, then we can apply the presented algorithms on every nontrivial elementary component of $G$.
Acknowledgment: Simon Brezovnik, Niko Tratnik, and Petra Žigert Pleteršek acknowledge the financial support from the Slovenian Research and Innovation Agency: research programme No. P1-0297 (Simon Brezovnik, Niko Tratnik, Petra Žigert Pleteršek), projects No. J1-4031 (Simon Brezovnik), N1-0285 (Niko Tratnik), and L7-4494 (Petra Žigert Pleteršek). All four authors thank the Slovenian Research and Innovation Agency for financing our bilateral project between Slovenia and the USA (title: Structural properties of resonance graphs and related concepts, project No. BI-US/22-24-158).